OpenStudy (anonymous):

Using the function g(x) = -2x + 6, compare and contrast g3) and g(-4). Which of the following statements is true?

OpenStudy (gorv):

g(3) means x=3 g(3)=-2*3+6=-6+6=0 similarly find g(-4) just replace x by -4